Holtzman Clothiers's stock currently sells for $24.00 a share. It just paid a dividend of $1.50 a share (i.e., D0 = $1.50). The dividend is expected to grow at a constant rate of 4% a year.

What stock price is expected 1 year from now? Round your answer to the nearest cent.
$  

What is the required rate of return? Do not round intermediate calculations. Round your answer to two decimal places.
  %
Answer #1

Answer:

D0 = $1.50

P0 = $24

G = 4%

D1 = D0 * (1+g)
D1 = $1.50 * (1 + 0.04)
D1 = $1.56

P1 = P0 * (1+g)
P1 = $24 * (1+0.04)
P1 = $24.96

R = D1 / P0 + g
R = $1.56 / $24 + 0.04
R = 0.065 + 0.04
R = 0.105 or 10.5%